OpenStudy (anonymous):

How many solutions does the equation have? 2(2x + 5 = 4(x + 3) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

A one solution

OpenStudy (anonymous):

the solution is -2

OpenStudy (anonymous):

Thanks alot! :)

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

is this the question ? : 2(2x + 5) = 4(x + 3)

OpenStudy (anonymous):

Yes but he told me the answer

OpenStudy (anonymous):

but there is no solution.

OpenStudy (anonymous):

2(2x+5)=4(x+3) 4x+10=4x+12 NO SOLUTION!!!

OpenStudy (anonymous):

@Kyliejones

OpenStudy (anonymous):

10 = 12 -> no solution

OpenStudy (anonymous):

So its no solution?

OpenStudy (anonymous):

yeah, no solution

OpenStudy (anonymous):

yes but please look at @Fallontine explanation

OpenStudy (anonymous):

Oh ok thank you guys!